Tom, Dick, and Harry weigh 50, 85, 120 lbs. respectively and use a 12ft plank for a teeterboard. Tom sits 5ft, Dick sits 4 ft, and Harry sits 3 ft from the fulcrum on the same side. The fulcrum is at the midpoint of the board. How far must a 200 lbs friend sit from the fulcrum to balance them.
Answer by checkley71(8403)

50*5=250
85*4=340
120*3=360
TOTAL FORCE=250+340+360=850.
THUS YOU HAVE 850 POUNDS OF FORCE TO OVERCOME.
A 200 POUND PERSON SITTING 850/200=4.25 FEET FROM THE FULCRUM WILL DO THE BALANCING ACT.